# JSJ splittings for all Artin groups

### v2

@@ Abstract.
We prove that an Artin group splits over infinite cyclic subgroups if and only if its defining graph has a separating vertex, and explicitly construct a JSJ decomposition over infinite cyclic subgroups for all Artin groups. This is then used to show that, if two Artin groups are isomorphic, then they have the same set of big chunks, that is, maximal subgraphs without separating vertices. We also deduce acylindrical hyperbolicity for the automorphism groups of many Artin groups, partially answering a question of Genevois in the case of Artin groups. As a consequence, this produces new families of Artin groups with the $R_{\infty}$ property.
@@ Introduction
Isomorphism invariance of big chunks.
One of the most widely open questions for Artin groups is the isomorphism problem, which asks for an algorithm which takes two labelled graphs $\Gamma$ and $\Gamma^{\prime}$ as input, and determines if they yield isomorphic Artin groups. This question, which specialises Dehn’s Isomorphism Problem to the family of Artin groups, is solved when one restricts to certain subclasses, including RAAGs [Dro87, Bau81], spherical-type [Par04], and large-type [Vas22], but there is no unified approach. However, isomorphic Artin groups must admit the same splittings over $\mathbb{Z}$, as this is an algebraic property. In this spirit, if $A_{\Gamma}\cong A_{\Gamma^{\prime}}$, then one can hope to use Theorem A to recover some decompositions of $\Gamma$ and $\Gamma^{\prime}$ into maximal connected subgraphs without separating vertices, which we call big chunks. We do just that in Theorem 5.6, of which we give a streamlined version here. For a labelled graph $\Gamma$, let $\mathcal{BC}(\Gamma)$ be the collection of its big chunks.
@@
Let $\Gamma$ and $\Gamma^{\prime}$ be finite, connected, labelled simplicial graphs, and let $\varphi\colon A_{\Gamma}\to A_{\Gamma^{\prime}}$ be an isomorphism. Then there exists a bijection $\varphi_{\#}\colon\mathcal{BC}(\Gamma)\to\mathcal{BC}(\Gamma^{\prime})$, such that for every $\Lambda\in\mathcal{BC}(\Gamma)$ the following hold:
@@
In other words, isomorphic Artin groups must have isomorphic big chunks, and, among these, the same number of even leaves of every given label. Theorem B is already new for two-dimensional Artin groups, and we believe that our findings, together with [Vas22, Theorem E], could be a key building block for a future solution of the isomorphism problem in this class.
@@
The key tool in the proof of Theorem B is the existence of a JSJ decomposition over cyclic subgroups, in the sense of [GL17]. We refer to Definition 3.2 for the details; for our purposes, it is enough to recall that, given a graph of group decomposition $\mathcal{G}$ of a group $G$, with cyclic edge groups, if $\mathcal{G}$-vertex groups are elliptic in every $G$-action on a tree with cyclic edge stabilisers, then $\mathcal{G}$ is a JSJ decomposition. In other words, a JSJ decomposition should be thought of as a “maximal” splitting of $G$, which means that vertex groups cannot be further split “in a canonical way”. Our methods allow us to build a JSJ decomposition over cyclic subgroups of an Artin group $A_{\Gamma}$, whose vertex groups roughly correspond to the big chunks of $\Gamma$:
@@
The construction of $J(\Gamma)$ is explicit, but we postpone it to Definition 3.7 and only provide an example here (see Figure 1). Crucially, the proof of Theorem 3.9 uses that big chunks on at least three vertices do not split further over $\mathbb{Z}$, as a consequence of Theorem A.
@@
Consider a connected graph $\Gamma$ as in Figure 1, whose big chunks are $\Sigma$, $\Lambda$, and the three leaves, with labels $2$, $3$, and $6$, respectively. The associated JSJ decomposition $J(\Gamma)$ of $A_{\Gamma}$ is also depicted in the Figure. White vertices correspond to separating vertices of $\Gamma$, while black vertices roughly correspond to big chunks. The only exception are even leaves: a leaf with label $2$ (here, spanned by $b$ and $v$) further splits as a HNN extension, while an even leaf with higher label (here, generated by $f$ and $g$) splits as an amalgamation $\mathbb{Z}^{2}*_{\mathbb{Z}}\mathbb{Z}$ over its centre (here, generated by $(fg)^{3}$, as the label is $6$). Whenever an edge of $J(\Gamma)$ is not labelled in this picture, the edge group is the white vertex group, whose embedding in the black vertex group corresponds to the graph inclusion of a separating vertex inside a big chunk.
@@
The cost of the tree of cylinders is that the edge groups are no longer cyclic. If one wishes to consider cyclic splittings, the canonical object is the JSJ deformation space. Deformation spaces were first introduced by Forester [For01], motivated by outer space for free groups [CV86]. Any JSJ decomposition of a group $G$ gives rise to the JSJ space [GL17], a contractible space with an action of the (outer) automorphism group of $G$. As such, the JSJ deformation space turns out to be a powerful tool to study (outer) automorphism groups; indeed, in [JMS25], we combine the results of this paper with the theory of deformation spaces to tackle a version of the isomorphism problem for new classes of Artin groups. More precisely, we are able to reduce the so-called Strong Twist Conjecture for an Artin group to its big chunks.
@@
Finally, the JSJ plays a central role in Section 5, where we prove that big chunks are isomorphism invariant, Theorem B (see Theorem 5.6).
@@ Acknowledgements
We thank Yassine Guerch and Gilbert Levitt for suggesting the application to acylindrical hyperbolicity of automorphism groups, and Francesco Fournier-Facio for pointing us towards its corollaries. We are also grateful to Kaitlin Ragosta and Ruta Sliazkaite for insightful discussions. A special thanks goes to our PhD supervisors Laura Ciobanu, Alessandro Sisto, and Alexandre Martin for comments on a first draft of this document.
@@ 2. Cyclic splittings of Artin groups
$z_{ab}=\begin{cases}\Delta_{ab}&\mbox{ if $m$ is even};\\ \Delta_{ab}^{2}&\mbox{ if $m$ is odd},\end{cases}$
@@
$g_{i+1}=\begin{cases}v_{j+1}&\mbox{ if }g_{i}=v_{j}\mbox{ and }m_{j}=2;\\ z_{j}&\mbox{ if }g_{i}=v_{j}\mbox{ and }m_{j}\geq 3;\\ v_{j+1}&\mbox{ if }g_{i}=z_{j}.\end{cases}$
@@
Case I. Assume first that each $G_{i}$ fixes a point. If some $G_{i}$ fixes an edge we are done; thus assume that each $G_{i}$ fixes a unique point $p_{i}$. If all the $p_{i}$ were the same, then $A_{\Gamma}$, which is generated by the $G_{i}$s, would have a global fixed point. So let $p$ be a leaf of the subtree $S$ spanned by the $p_{i}$, and let $I\subseteq\{1,\ldots,r\}$ be such that $G_{i}$ fixes $p$ if and only if $i\in I$. Notice that $I$ is non-empty, and does not contain all indexes as there is no global fixed point; hence there exist $j,k\in I$ such that $j-1,k+1\not\in P$, where addition is modulo $r$. Let $p_{j-1}$ (resp. $p_{k+1}$) be the point fixed by $G_{j-1}$ (resp. $G_{k+1}$). This way, $g_{j}\in G_{j-1}\cap G_{j}$ fixes the non-trivial segment $[p_{j-1},p]$, and similarly $g_{k+1}\in G_{k}\cap G_{k+1}$ fixes $[p_{k+1},p]$. As $p$ was a leaf of $S$, the two segments share an edge, so $\langle g_{j},g_{k+1}\rangle$ fixes an edge. Now, $g_{i}$ and $g_{k+1}$ are infinite order elements without a non-trivial common power, as they are either two different vertices, or a vertex and the generator of a centre, or the generators of the centres of two different dihedrals (in all three cases, the absence of a common power follows immediately from [VdL83]). In particular $\langle g_{j},g_{k+1}\rangle$ is not virtually cyclic and fixes an edge, proving that $A_{\Gamma}$ does not split over virtually cyclic groups.
@@
Since there is no global fixed point, there exist vertices $v,v^{\prime}$ of $\Gamma$ which do not share a fixed point. As argued above, $v$ and $v^{\prime}$ cannot belong to the same two-edges segment, so they are at distance $n\geq 3$ in $\Gamma$. Furthermore, if we take a combinatorial path $v=v_{0},\ldots,v_{n}=v^{\prime}$ from $v$ to $v^{\prime}$, there must be some index $10<i<n-$ 1 such that $\langle v_{i-1},v_{i},v_{i+1}\rangle$ and $\langle v_{i},v_{i+1},v_{i+2}\rangle$ fix different points. In particular, the dihedral $\langle v_{i},v_{i+1}\rangle$ fixes an edge, so $A_{\Gamma}$ does not split over virtually cyclic groups. ∎
@@ A $VC$-splitting $(T,\Omega)$ of a group $G$ is a JSJ tree over virtually cyclic subgroups if:
Consider the bipartite graph $\mathcal{B}(\Gamma)$ defined as follows:
@@ 4.3. Application to automorphism groups of Artin groups
5. Isomorphism invariance of big chunks
Using the JSJ decompositions we constructed, in this final Section we prove that, if two Artin groups are isomorphic, then any isomorphism must preserve the isomorphism type of big chunks. In fact, we show that every isomorphism preserves the conjugacy class of any big chunk that is not a toral leaf.
@@ 5. Isomorphism invariance of big chunks
Theorem 5.6 (Big chunks are isomorphism invariants).
@@
To see that $\varphi(\langle a,b\rangle)\leq k\langle c,d\rangle k^{-1}$, let $\psi$ be the composition of $\varphi$ and the conjugation by $k^{-1}$, so that we have to show that $\psi(\langle a,b\rangle)\leq\langle c,d\rangle$. In turn, it is enough to prove that $\psi(r)\leq\langle c,d\rangle$ where $r=ab$. Since $r^{m}=z_{ab}$, the element $x\coloneq\psi(r)^{m}$ is primitive inside $\langle c,z_{cd}\rangle$; in particular, $\psi(r)$ is elliptic and fixes some coset $gW$, where $g\in A_{\Gamma^{\prime}}$ and $W$ is a vertex group of $J(\Gamma^{\prime})$. Let $\mathcal{T}$ be the sub-tree of $J(\Gamma^{\prime})$ corresponding to the $\langle c,d\rangle$-orbits of the edge between the black vertex $\langle c,z_{cd}\rangle$ and the red vertex $\langle cd\rangle$. If $gW\in\mathcal{T}$ we are done, because then both $g$ and $W$ belong to $\langle c,d\rangle$ and so $\psi(r)\in gWg^{-1}\leq\langle c,d\rangle$. Otherwise, $x$ fixes the geodesic from $gW$ to $\langle c,z_{cd}\rangle$, and on this geodesic let $e$ be the last edge not contained in $\mathcal{T}$. By how $J(\Gamma^{\prime})$ is constructed, we must have that $e=h\langle c\rangle$, where $h$ lies in some $\langle c,d\rangle$-translate of $\langle c,z_{cd}\rangle$, and in particular in $\langle c,d\rangle$. Then $x\in h\langle c\rangle h^{-1}\cap\langle c,z_{cd}\rangle$. Since no power of $hch^{-1}$ is central in $\langle c,d\rangle$, the image $\overline{x}$ of $x$ in the central quotient $\langle c,d\rangle/\langle z_{cd}\rangle\cong\mathbb{Z}*\mathbb{Z}/m\mathbb{Z}$ is non-trivial, and lies in the intersection between $\langle\overline{c}\rangle$ and $\langle\overline{h}\overline{c}\overline{h}^{-1}\rangle$. Since $\langle\overline{c}\rangle$ generates the $\mathbb{Z}$-factor in the free product, which is malnormal, we have that $\overline{h}\in\langle\overline{c}\rangle$, so $h\in\langle c,z_{cd}\rangle$ and therefore $hch^{-1}=c$. In turn, $x\in\langle c\rangle$ is primitive, so $x=c^{\pm 1}$; however, sending every standard generator to $1$ defines a map $A_{\Gamma^{\prime}}\to\mathbb{Z}$, which maps $\psi(r)$ to an $m$-th root of ±$\pm 1$, a contradiction.

### v3

@@ Abstract.
We prove that an Artin group splits over infinite cyclic subgroups if and only if its defining graph has a separating vertex, and explicitly construct a JSJ decomposition over infinite cyclic subgroups for all Artin groups. We then use these facts to show that, if two Artin groups are isomorphic, then they have the same set of parabolics supported on big chunks, that is, maximal subgraphs without separating vertices. We also deduce acylindrical hyperbolicity for the automorphism groups of many Artin groups, partially answering a question of Genevois in the case of Artin groups. As a consequence, we produce new families of Artin groups with the $R_{\infty}$ property.
@@ Introduction
Isomorphism invariance of big chunk parabolics.
One of the most widely open questions for Artin groups is the isomorphism problem, which asks for an algorithm which takes two labelled graphs $\Gamma$ and $\Gamma^{\prime}$ as input, and determines if they yield isomorphic Artin groups. This question, which specialises Dehn’s Isomorphism Problem to the family of Artin groups, is solved when one restricts to certain subclasses, including RAAGs [Dro87, Bau81], spherical-type [Par04], and large-type [Vas22], but there is no unified approach. However, isomorphic Artin groups must admit the same splittings over $\mathbb{Z}$, as this is an algebraic property. In the same spirit, if two Artin groups are isomorphic, one can hope to use Theorem A to recover a correspondence between “maximal” parabolic subgroups which do not split over $\mathbb{Z}$.
To make the above more precise, we first need a definition. Given a labelled graph $\Gamma$, a big chunk is a connected induced subgraph without separating vertices, which is maximal with these properties. [^1] For unlabelled simplicial graphs, a big chunk is also called a biconnected component or a block. Though the latter terminology is more widespread, especially in computer science, we preferred the former, as it was recently used in the literature in the context of Artin groups. In Theorem 5.6, of which we give a streamlined version here, we prove that isomorphic Artin groups have isomorphic big chunk parabolics, that is, parabolic subgroups supported on big chunks:
@@
Let $\Gamma$ and $\Gamma^{\prime}$ be finite, connected, labelled simplicial graphs, and let $\mathcal{BC}(\Gamma)$ and $\mathcal{BC}(\Gamma^{\prime})$ be the collection of big chunks of $\Gamma$ and $\Gamma^{\prime}$, respectively. Furthermore, let $\varphi\colon A_{\Gamma}\to A_{\Gamma^{\prime}}$ be an isomorphism. Then there exists a bijection $\varphi_{\#}\colon\mathcal{BC}(\Gamma)\to\mathcal{BC}(\Gamma^{\prime})$, such that for every $\Lambda\in\mathcal{BC}(\Gamma)$ the following hold:
@@
Theorem B is already new for two-dimensional Artin groups, and we believe that our findings, together with [Vas22, Theorem E], could be a key building block for a future solution of the isomorphism problem in this class.
@@
The key tool in the proof of Theorem B is the existence of a JSJ decomposition over cyclic subgroups, in the sense of [GL17]. We refer to Definition 3.2 for the details; for our purposes, it is enough to recall that, given a graph of group decomposition $\mathcal{G}$ of a group $G$, with cyclic edge groups, if $\mathcal{G}$-vertex groups are elliptic in every $G$-action on a tree with cyclic edge stabilisers, then $\mathcal{G}$ is a JSJ decomposition. In other words, a JSJ decomposition should be thought of as a “maximal” splitting of $G$, which means that vertex groups cannot be further split “in a canonical way”. Our methods allow us to build a JSJ decomposition over cyclic subgroups of an Artin group $A_{\Gamma}$, whose vertex groups roughly correspond to the big chunk parabolics of $\Gamma$:
@@
The construction of $J(\Gamma)$ is explicit, but we postpone it to Definition 3.7 and only provide an example here (see Figure 1). Crucially, the proof of Theorem 3.9 uses that big chunk parabolics on at least three vertices do not split further over $\mathbb{Z}$, as a consequence of Theorem A.
@@
Consider a connected graph $\Gamma$ as in Figure 1, whose big chunks are $\Sigma$, $\Lambda$, and the three leaves, with labels $2$, 33, and $6$, respectively. The associated JSJ decomposition $J(\Gamma)$ of $A_{\Gamma}$ is also depicted in the Figure. White vertices correspond to separating vertices of $\Gamma$, while black vertices roughly correspond to big chunks. The only exception are even leaves: a leaf with label $2$ (here, spanned by $b$ and $v$) further splits as a HNN extension, while an even leaf with higher label (here, generated by $f$ and $g$) splits as an amalgamation $\mathbb{Z}^{2}*_{\mathbb{Z}}\mathbb{Z}$ over its centre (here, generated by $(fg)^{3}$, as the label is $6$). Whenever an edge of $J(\Gamma)$ is not labelled in this picture, the edge group is the white vertex group, whose embedding in the black vertex group corresponds to the graph inclusion of a separating vertex inside a big chunk.
@@
The cost of the tree of cylinders is that the edge groups are no longer cyclic. If one wishes to consider cyclic splittings, the canonical object is the JSJ deformation space. Deformation spaces were first introduced by Forester [For01], motivated by outer space for free groups [CV86]. Any JSJ decomposition of a group $G$ gives rise to the JSJ space [GL17], a contractible space with an action of the (outer) automorphism group of $G$. As such, the JSJ deformation space turns out to be a powerful tool to study (outer) automorphism groups; indeed, in [JMS25], we combine the results of this paper with the theory of deformation spaces to tackle a version of the isomorphism problem for new classes of Artin groups. More precisely, we are able to reduce the so-called Strong Twist Conjecture for an Artin group to its big chunk parabolics.
@@
Finally, the JSJ plays a central role in Section 5, where we prove that big chunk parabolics are isomorphism invariant, Theorem B (see Theorem 5.6).
@@ Acknowledgements
We thank Yassine Guerch and Gilbert Levitt for suggesting the application to acylindrical hyperbolicity of automorphism groups, and Francesco Fournier-Facio for pointing us towards its corollaries. We are also grateful to Kaitlin Ragosta for enlightenments on spherical Artin groups, to Bianca Marchionna for pointing us towards the connections to graph theory, and to Ruta Sliazkaite for insightful discussions. A special thanks goes to our PhD supervisors Laura Ciobanu, Alessandro Sisto, and Alexandre Martin for comments on a first draft of this document.
@@ 2. Cyclic splittings of Artin groups
[…]
@@
[…]
@@
Case I. Assume first that each $G_{i}$ fixes a point. If some $G_{i}$ fixes an edge we are done; thus assume that each $G_{i}$ fixes a unique point $p_{i}$. If all the $p_{i}$ were the same, then $A_{\Gamma}$, which is generated by the $G_{i}$s, would have a global fixed point. So let $p$ be a leaf of the subtree $S$ spanned by the $p_{i}$, and let $I\subseteq\{1,\ldots,r\}$ be such that $G_{i}$ fixes $p$ if and only if $i\in I$. Notice that $I$ is non-empty, and does not contain all indexes as there is no global fixed point; hence there exist $j,k\in I$ such that $j-1,k+1\not\in P$, where addition is modulo $r$. Let $p_{j-1}$ (resp. $p_{k+1}$) be the point fixed by $G_{j-1}$ (resp. $G_{k+1}$). This way, $g_{j}\in G_{j-1}\cap G_{j}$ fixes the non-trivial segment $p_{j-1},p]$, and similarly $g_{k+1}\in G_{k}\cap G_{k+1}$ fixes $p_{k+1},p]$. As $p$ was a leaf of $S$, the two segments share an edge, so $\langle g_{j},g_{k+1}\rangle$ fixes an edge. Now, $g_{i}$ and $g_{k+1}$ are infinite order elements without a non-trivial common power, as they are either two different vertices, or a vertex and the generator of a centre, or the generators of the centres of two different dihedrals (in all three cases, the absence of a common power follows immediately from [VdL83]). In particular $\langle g_{j},g_{k+1}\rangle$ is not virtually cyclic and fixes an edge, proving that $A_{\Gamma}$ does not split over virtually cyclic groups.
@@
Since there is no global fixed point, there exist vertices $v,v^{\prime}$ of $\Gamma$ which do not share a fixed point. As argued above, $v$ and $v^{\prime}$ cannot belong to the same two-edges segment, so they are at distance $n\geq 3$ in $\Gamma$. Furthermore, if we take a combinatorial path $v=v_{0},\ldots,v_{n}=v^{\prime}$ from $v$ to $v^{\prime}$, there must be some index 0<i<n10<i<n-1 such that $\langle v_{i-1},v_{i},v_{i+1}\rangle$ and $\langle v_{i},v_{i+1},v_{i+2}\rangle$ fix different points. In particular, the dihedral $\langle v_{i},v_{i+1}\rangle$ fixes an edge, so $A_{\Gamma}$ does not split over virtually cyclic groups. ∎
@@ A $VC$-splitting $(T,\Omega)$ of a group $G$ is a JSJ tree over virtually cyclic subgroups if:
Let $\Gamma$ be a finite, labelled, simplicial graph, and let $\mathcal{B}(\Gamma)$ be the bipartite graph defined as follows: [^2] In graph theory, $\mathcal{B}(\Gamma)$ is known as the block-cut tree of $\Gamma$.
@@ 4.3. Application to automorphism groups of Artin groups
5. Isomorphism invariance of big chunk parabolics
Using the JSJ decompositions we constructed, in this final Section we prove that, if two Artin groups are isomorphic, then any isomorphism must preserve the isomorphism type of big chunk parabolics. In fact, we show that every isomorphism preserves the conjugacy class of any big chunk parabolic that is not a toral leaf.
@@ 5. Isomorphism invariance of big chunk parabolics
Theorem 5.6 (Big chunk parabolics are isomorphism invariants).
@@
To see that $\varphi(\langle a,b\rangle)\leq k\langle c,d\rangle k^{-1}$, let $\psi$ be the composition of $\varphi$ and the conjugation by $k^{-1}$, so that we have to show that $\psi(\langle a,b\rangle)\leq\langle c,d\rangle$. In turn, it is enough to prove that $\psi(r)\leq\langle c,d\rangle$ where $r=ab$. Since $r^{m}=z_{ab}$, the element $x\coloneq\psi(r)^{m}$ is primitive inside $\langle c,z_{cd}\rangle$; in particular, $\psi(r)$ is elliptic and fixes some coset $gW$, where $g\in A_{\Gamma^{\prime}}$ and $W$ is a vertex group of $J(\Gamma^{\prime})$. Let $\mathcal{T}$ be the sub-tree of $J(\Gamma^{\prime})$ corresponding to the $\langle c,d\rangle$-orbits of the edge between the black vertex $\langle c,z_{cd}\rangle$ and the red vertex $\langle cd\rangle$. If $gW\in\mathcal{T}$ we are done, because then both $g$ and $W$ belong to $\langle c,d\rangle$ and so $\psi(r)\in gWg^{-1}\leq\langle c,d\rangle$. Otherwise, $x$ fixes the geodesic from $gW$ to $\langle c,z_{cd}\rangle$, and on this geodesic let $e$ be the last edge not contained in $\mathcal{T}$. By how $J(\Gamma^{\prime})$ is constructed, we must have that $e=h\langle c\rangle$, where $h$ lies in some $\langle c,d\rangle$-translate of $\langle c,z_{cd}\rangle$, and in particular in $\langle c,d\rangle$. Then $x\in h\langle c\rangle h^{-1}\cap\langle c,z_{cd}\rangle$. Since no power of $hch^{-1}$ is central in $\langle c,d\rangle$, the image $\overline{x}$ of $x$ in the central quotient $\langle c,d\rangle/\langle z_{cd}\rangle\cong\mathbb{Z}*\mathbb{Z}/m\mathbb{Z}$ is non-trivial, and lies in the intersection between $\langle\overline{c}\rangle$ and $\langle\overline{h}\overline{c}\overline{h}^{-1}\rangle$. Since $\langle\overline{c}\rangle$ generates the $\mathbb{Z}$-factor in the free product, which is malnormal, we have that $\overline{h}\in\langle\overline{c}\rangle$, so $h\in\langle c,z_{cd}\rangle$ and therefore $hch^{-1}=c$. In turn, $x\in\langle c\rangle$ is primitive, so $x=c^{\pm 1}$; however, sending every standard generator to $1$ defines a map $A_{\Gamma^{\prime}}\to\mathbb{Z}$, which maps $\psi(r)$ to an $m$-th root of $\pm 1$, a contradiction.